Mike deposited $500 for 9months at 8%. Compounded quarterly
A) How many times was interest added to Mike's account?
B) What percent intrest was added each time?
C) What was the balance in Mike's account at the end of one year?

Answer by stanbon(75887)

Mike deposited $500 for 9months at 8%. Compounded quarterly
A(t) = P(1+(r/n))^(nt)
-------------------------------------
A(3/4) = 500(1+(0.08/4))^(4*(3/4))
------
A) How many times was interest added to Mike's account?
Ans: 4*(3/4) = 3
---------------------------
B) What percent intrest was added each time?
Ans: 0.08/4 = 0.02 = 2%
-----------------------------------
C) What was the balance in Mike's account at the end of one year?
A(1) = 500(1+(0.08/4))^(4*1)
A(1) = 500(1.02)^4
A(1) = $541.22
============================
Cheers,
Stan H.
============================